Please help me on this problem:
Solve the linear system using the substitution method.
-4x+3y=-5
x=4-2y
thank you

Answer by boilpoil(127)

-4x+3y=-5
x=4-2y ---Substitute the subject x into first equation

-4(4-2y)+3y=-5
-16+8y+3y=-5
11y=11
y=1

Value of x
=4-2(1)
=4-2
=2